Direct Labor Variances

The following data relate to labor cost for production of 3,800 cellular telephones:

Actual: 2,560 hrs. at $12.9
Standard: 2,520 hrs. at $13.1

a. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Rate variance $   
Time variance $   
Total direct labor cost variance $   

b. The employees may have been less-experienced or poorly trained, thereby resulting in a   labor rate than planned. The lower level of experience or training may have resulted in   efficient performance. Thus, the actual time required was   than standard.
Answer #1

a. Rate variance = (standard rate - actual rate)*actual hours

= (13.1-12.9)2560 = $512 favorable

Time variance = (standard hours-actual hours)*standard rate

= (2520-2560)13.1 = $524 unfavorable

Total direct labor cost variance = standard cost - actual cost

= 2520*13.1-2560*12.9 = $12 unfavorable

b. The employees may have been less-experienced or poorly trained, thereby resulting in a lower labor rate than planned. The lower level of experience or training may have resulted in less efficient performance. Thus, the actual time required was higher than standard.